# Foundation Model Priors Enhance Object Focus in Feature Space for Source-Free Object Detection

### v1

@@ 1 Introduction
Figure 1: Towards Object-Focused Representations in SFOD. Two examples from the Foggy Cityscapes [sakaridis2018semantic] target set. In each example, the two columns show the student’s thresholded channel-mean feature maps from the last backbone layer (brighter = higher activation) and the model predictions, respectively. Existing state-of-the-art methods IRG [irg] and Simple-SFOD [hao2024simplifying] tend to produce less object-focused activations extending into background regions, leading to less precise object localization (red arrows) and false positives (orange arrows). In contrast, our method produces compact, object-shaped activations with tighter boxes and accurate labels, demonstrating stronger spatial coherence and object awareness. (Best viewed when zoomed in)
@@
Source-Free Object Detection (SFOD) aims to adapt a detector trained on labeled source data to an unlabeled target domain without accessing any source samples during adaptation. This setting is practically important, as it removes the need to store or share potentially sensitive source data, a common constraint in real-world applications such as autonomous driving, surveillance, and medical imaging. Current state-of-the-art SFOD approaches [irg, pets, hao2024simplifying, khanh2024dynamic] largely adopt the Mean Teacher self-labeling framework. In this paradigm, a teacher network, maintained as an exponential moving average (EMA) of the student, generates pseudo-labels on target images to facilitate adaptation. Although this framework effectively exploits unlabeled data, it inherits a strong source-domain bias. Since the teacher is trained exclusively on source data, it often produces unreliable pseudo-labels when applied to the target-domain [sfod, sfod_mosaic, a2sfod, lods, irg, pets, hao2024simplifying, khanh2024dynamic]. These noisy labels are subsequently propagated to the student, thereby compromising adaptation stability. To alleviate this issue, existing methods refine pseudo-label selection [sfod, sfod_mosaic, pets, khanh2024dynamic], exploit object relationships within the target-domain [irg], enforce style invariance [lods], or apply adversarial alignment strategies [a2sfod]. However, these approaches largely overlook a more fundamental limitation. We observe that domain shift weakens the detector’s object focus, wherein feature activations become spatially diffuse and extend into background clutter, reducing the clarity of object boundaries, as demonstrated in Figure 1. As a result, the detection head operates on less discriminative features, producing suboptimal pseudo-labels. This observation motivates us to strengthen SFOD from the feature level by promoting structured, object-centered representations rather than solely refining pseudo-labels.
@@
Recent progress in cross-domain object detection [lavoie2025large, han2025vfm] demonstrates that large-scale vision foundation models can serve as reliable, domain-agnostic knowledge sources. Motivated by these advances, we explore their potential in a source-free setting and propose SPAR (Spatial Prior-Aware Regularization), a framework that uses class-agnostic binary masks derived from a frozen segmentation model [yuan2024open] to regularize the detector’s feature space. SPAR encourages the student’s channel-mean activations to align with these masks, enhancing foreground structure and promoting object-focused responses. While SPAR improves the spatial quality of learned features, pseudo-labels in detection remain highly imbalanced, with abundant background regions and scarce positive samples. To stabilize training under this imbalance and mitigate residual class-label noise, we introduce IRPL (Imbalance-aware Noise Robust Pseudo-Labeling). It is a classification loss designed to promote balanced and noise-tolerant learning by down-weighting overconfident teacher-student agreements, reinforcing underrepresented categories, and countering the inherent background dominance in detection. Together, SPAR and IRPL complement each other: SPAR strengthens the structural integrity of object representations, while IRPL ensures balanced and robust supervision. We further provide a theoretical analysis linking these contributions to tighter localization and classification error bounds. Overall, our framework advances object-focused adaptation under domain shift. Our main contributions are summarized as follows:
@@ 2 Related Works
UDAOD adapts a source-trained detector to an unlabeled target-domain when source data is available during adaptation [robust, mega_cdn, munir2021ssal, kennerley2024cat, li2025seen]. Recent works like DINO Teacher [lavoie2025large] leverage foundation models by training a source-domain labeller with a frozen DINOv2 and then aligning the student’s patch features to a frozen DINO encoder. While this approach enriches supervision, it remains computationally demanding as the DINO encoder is used in an online manner for target adaptation. Our SPAR loss provides a contrasting design: it uses class-agnostic binary masks from a frozen open-vocabulary segmenter [yuan2024open] that are computed once before adaptation, so the foundation model is not queried during training or inference and the added cost is a one-time preprocessing step.
@@
SFOD eliminates the need to access source data during adaptation [irg, pets, hao2024simplifying, varailhon2024source, khanh2024dynamic]. IRG [irg] refines pseudo-labels via instance-relation graphs, PETS [pets] stabilizes teacher-student training with periodic exchanges, and Simple-SFOD [hao2024simplifying] shows that careful self-training design can outperform complex architectures. SF-YOLO [varailhon2024source] adapts the YOLO detector in a source-free manner using target domain-specific augmentations and a teacher-student communication mechanism. DRU [khanh2024dynamic] dynamically retrains and updates the mean teacher to enhance stability and performance during adaptation. While these works focus on refining pseudo-labels or stabilizing training, they overlook the importance of object-focused feature representation. Our proposed model addresses both the problems in a lightweight (simple-by-design, zero inference overhead), effective manner.
@@
Noisy-label learning introduces objectives that suppress the impact of corrupted labels [wang2024epsilon]. However, a direct adaptation of  [wang2024epsilon] fails in SFOD (detailed in Section 5.2) due to proposal imbalance, heavy class skew, and confirmation bias in teacher-student agreements. We extend it with foreground/background weighting, and entropy regularization, which together make IRPL effective for detection.
@@ 3.2 FALCON-SFOD
Motivation and Overview. As observed in [irg, pets, hao2024simplifying, khanh2024dynamic], a prevalent problem in mean-teacher self-training is that pseudo class-labels tend to be noisy due to domain shifts. Beyond this well-known issue, we identify that domain shift weakens the detector’s object-focus in SFOD: spatial activations for true objects are diluted by background clutter, which degrades localization and cascades into misclassification (see Fig. 1). We propose FALCON-SFOD, coupling two dedicated objectives within the standard Mean-Teacher framework: SPAR for spatial focus regularization and IRPL for pseudo class-label robustness. These losses are integrated with the conventional localization objective to enhance robustness during source-free adaptation.
Spatial Prior-Aware Regularization (SPAR). Consider the detector $h(x)=f(g(x))$ (Sec. 3.1). The feature extractor outputs an activation tensor $a=g(x)\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ are its height, width, and channel count. Taking the channel-wise mean yields $A\in\mathbb{R}^{H\times W}$, which highlights spatial locations with high average activations and thus usually traces foreground objects [rebbapragada2024c2fdrone]. Under a domain shift, however, this map becomes contaminated by background clutter: true objects may fade while background areas are spuriously accentuated (as illustrated in Fig. 1). We derive a foreground prior once, offline, by running a frozen open-vocabulary segmentation model on target images. We discard the class information and binarize these masks, where pixels belonging to any segmented region are set to $1$ (foreground) and the rest to 0 (background). The prior is not queried during training or inference. For a target image $x_{i}^{t}$, let $A_{G}(x_{i}^{t})\in[0,1]^{H^{\prime}\times W^{\prime}}$ be the prior mask and $A_{S}(x_{i}^{t})\in[0,1]^{H\times W}$ be the student channel-mean map rescaled to $H^{\prime}\times W^{\prime}$ to match the prior. We encourage agreement using a mean $\ell_{1}$ term plus Dice::
@@
Imbalance-aware Noise Robust Pseudo-Labeling (IRPL). Most state-of-the-art SFOD methods minimize cross-entropy loss [irg, pets, hao2024simplifying], which has an unbounded gradient that allows a single corrupted pseudo class-label to dominate training. Additionally, object detection has an inherent foreground-background imbalance which affects the stability of the student training. We therefore introduce IRPL, which mitigates this negative impact of noisy pseudo class-labels by adaptively recalibrating the per-box classification loss and effectively re-weights foreground-vs-background in the loss. For student probabilities $f_{c}^{st}\!\in\!\mathbb{R}^{K+1}$, let $\mathbf{p}=f_{c}^{st}$ and $t=\arg\max_{k}p_{k}$. The following transform rescales the peak while making sure that elements of $\mathbf{p}^{\prime}$ sum up to 1.
@@ 4 Theoretical Insights
Lemma 1 employs the class-conditional transition matrix $T\!\in\![0,1]^{(K+1)\times(K+1)}$, where $T_{ji}=\Pr[\hat{c}=i\mid c=j]$ represents the label noise induced by the mean-teacher: an exponential-moving-average (EMA) copy of the student whose deterministic predictions form a stochastic channel over the data distribution [mean_teacher]. The diagonal element $T_{jj}$ is the teacher’s per-class hit-rate, and we define $\lambda=\min_{j}T_{jj}>0$ to rule out classes the teacher never recognizes, a minimal condition for identifiability of the clean risk under arbitrary label noise [liu2023identifiability]. Lemma 1 therefore shows that mean-teacher asymmetry costs only a multiplicative factor $1/\lambda$; when the teacher is perfect ($\lambda=1$) the bound reduces to the standard clean-risk expression. Next, we show in Lemma 2 the regression risk under noisy pseudo bounding box labels.
@@
UDAOD SSAL [munir2021ssal] (NeurIPS’21) 45.1 47.4 59.4 24.5 50 25.7 26 38.7 39.6 51.8 45.6
PT [pt] (ICML’22) 40.2 48.8 59.7 30.7 51.8 30.6 35.4 44.5 42.7 55.1 60.2
MTM [weng2024mean] (AAAI’24) 51 53.4 67.2 37.2 54.4 41.6 38.4 47.7 48.9 58.1 -
SEEN-DA [li2025seen] (CVPR’25) 58.5 64.5 71.7 42 61.2 54.8 47.1 59.9 57.5 66.8 67.1
SFOD SFOD [sfod] (AAAI’21) 21.7 44.0 40.4 32.2 11.8 25.3 34.5 34.3 30.6 42.3 43.6
SFOD-Mosaic [sfod_mosaic] (AAAI’21) 25.5 44.5 40.7 33.2 22.2 28.4 34.1 39.0 33.5 42.9 44.6
LODS [lods] (CVPR’22) 34.0 45.7 48.8 27.3 39.7 19.6 33.2 37.8 35.8 - 43.9
IRG [irg] (CVPR’23) 37.4 45.2 51.9 24.4 39.6 25.2 31.5 41.6 37.1 45.2 46.9
PETS [pets] (ICCV’23) 42.0 48.7 56.3 19.3 39.3 5.5 34.2 41.6 35.9 57.8 47.0
Simple-SFOD [hao2024simplifying] (ECCV’24) 40.9 48 58.9 29.6 51.9 50.2 36.2 44.1 45.0 55.4 46.2
SF-YOLO [varailhon2024source] (ECCV’24 w) 47.4 50.6 64.1 26.0 49.7 32.5 25.9 44.1 42.5 57.7 49.4
DRU [khanh2024dynamic] (ECCV’24) 48.3 51.5 62.5 26.2 43.2 34.1 34.2 48.6 43.7 58.7 45.1
@@ 5 Experiments
Datasets and Metrics. We use five publicly available datasets covering four domain shift scenarios. Cityscapes [cordts2016cityscapes] is an urban street scene dataset comprising 5,000 finely annotated images collected from diverse cities and seasons, from which we use 2,925 images for training and 500 for validation. It includes eight categories: person, rider, car, truck, bus, train, motorcycle, and bicycle. Foggy Cityscapes [sakaridis2018semantic] extends Cityscapes by overlaying synthetic fog at three intensity levels (0.005, 0.01, and 0.02) to simulate poor visibility conditions. KITTI [geiger2013vision] is an autonomous driving dataset consisting of 7,481 real-world street scene training images. Sim10k [johnson2016driving] provides 10,000 synthetic urban scene images of cars, rendered from the video game Grand Theft Auto. BDD100k [yu2018bdd100k] is a large-scale dataset comprising 100,000 driving scene images captured across various weather conditions. Following existing works, we report mean average precision (mAP) at 0.5 IoU threshold. To demonstrate the efficacy of our method on extreme domain shifts, we use following datasets: PascalVOC [everingham2010pascal], COCO [lin2014microsoft], FLIR [flir2019adas], and Clipart [inoue2018cross].
Implementation Details. Our student model is trained using SGD with learning rates of 0.04 for source-only training and 0.0025 for target-domain adaptation, and a batch size of 4. The teacher model is updated via EMA with decay factor $\delta$ = 0.9996, using a confidence threshold 0.8. We use  [yuan2024open] for getting the class-agnostic binary foreground masks for SPAR loss. The training parameters are identical for every dataset. All experiments are run on one NVIDIA RTX A6000 GPU. Results in the paper use [hao2024simplifying] as the base model unless specified otherwise. Similar to  [irg, hao2024simplifying], we use 0.02 fog-level images from the Foggy Cityscapes for adaptation. Our code will be made public upon acceptance.
@@ 5.1 Comparisons with State-of-the-art
Adaptation to Adverse Weather. To evaluate the adverse weather domain shifts, we perform adaptation from Cityscapes to Foggy Cityscapes. Table 1 shows that our method achieves a state-of-the-art (sota) mAP of 46.9%, outperforming DRU [khanh2024dynamic] by $3.2\%$, SF-YOLO [varailhon2024source] by $4.9\%$, and Simple-SFOD [hao2024simplifying] by $1.9\%$. Especially, our method achieves notable improvements of 4.1%, 2.9%, and 2.4% in the under-represented and challenging classes train, bus, and motorcycle, respectively, compared to the previous sota in those categories.
@@
Small-scale to Large-scale. We select Cityscapes as the source-domain and BDD100k as the target-domain to study this shift. Following [pets, hao2024simplifying], we focus on the seven categories shared with Cityscapes. As shown in Table 2, our method achieves an mAP of 36.9%, beating Simple-SFOD [hao2024simplifying] by 2.6%.
@@
UDAOD SWDA [swda] (CVPR’19) 15.2 45.7 29.5 30.2 17.1 21.2 18.4 25.3
CR-DA-Det [cat_reg] (CVPR’20) 19.5 46.3 31.3 31.4 17.3 23.8 18.9 26.9
MTM [weng2024mean] (AAAI’24) 53.7 35.1 68.8 23.0 28.8 23.8 28.0 37.3
SFOD SFOD [sfod] (AAAI’21) 20.4 48.8 32.4 31.0 15.0 24.3 21.3 27.6
SFOD-Mosaic [sfod_mosaic] (AAAI’21) 20.6 50.4 32.6 32.4 18.9 25.0 23.4 29.0
A2SFOD [a2sfod] (AAAI’23) 26.6 50.2 36.3 33.2 22.5 28.2 24.4 31.6
IRG [irg] (CVPR’23) 31.4 59.7 32.8 39.9 16.7 26.9 21.5 32.7
PETS [pets] (ICCV’23) 19.3 62.4 34.5 42.6 17.0 26.3 16.9 31.3
Simple-SFOD [hao2024simplifying] (ECCV’24) 32 60 33.4 40.2 19.7 29.9 24.9 34.3
DRU [khanh2024dynamic] (ECCV’24) 27.1 62.7 36.9 45.8 22.7 32.5 28.1 36.6

### v3

@@ 1 Introduction
Figure 1: Towards Object-Focused Representations in SFOD. Two examples from the Foggy Cityscapes [26] target set. In each example, the two columns show the student’s thresholded channel-mean feature maps from the last backbone layer (brighter = higher activation) and the model predictions, respectively. Existing state-of-the-art methods IRG [30] and Simple-SFOD [7] tend to produce less object-focused activations extending into background regions, leading to less precise object localization (red arrows) and false positives (orange arrows). In contrast, our method produces compact, object-shaped activations with tighter boxes and accurate labels, demonstrating stronger spatial coherence and object awareness. (Best viewed when zoomed in)
@@
Source-Free Object Detection (SFOD) aims to adapt a detector trained on labeled source data to an unlabeled target domain without accessing any source samples during adaptation. This setting is practically important, as it removes the need to store or share potentially sensitive source data, a common constraint in real-world applications such as autonomous driving, surveillance, and medical imaging. Current state-of-the-art SFOD approaches [30, 20, 7, 11] largely adopt the Mean Teacher self-labeling framework. In this paradigm, a teacher network, maintained as an exponential moving average (EMA) of the student, generates pseudo-labels on target images to facilitate adaptation. Although this framework effectively exploits unlabeled data, it inherits a strong source-domain bias. Since the teacher is trained exclusively on source data, it often produces unreliable pseudo-labels when applied to the target-domain [17, 18, 2, 16, 30, 20, 7, 11]. These noisy labels are subsequently propagated to the student, thereby compromising adaptation stability. To alleviate this issue, existing methods refine pseudo-label selection [17, 18, 20, 11], exploit object relationships within the target-domain [30], enforce style invariance [16], or apply adversarial alignment strategies [2]. However, these approaches largely overlook a more fundamental limitation. We observe that domain shift weakens the detector’s object focus, wherein feature activations become spatially diffuse and extend into background clutter, reducing the clarity of object boundaries, as demonstrated in Figure 1. As a result, the detection head operates on less discriminative features, producing suboptimal pseudo-labels. This observation motivates us to strengthen SFOD from the feature level by promoting structured, object-centered representations rather than solely refining pseudo-labels.
@@
Recent progress in cross-domain object detection [13, 6] demonstrates that large-scale vision foundation models can serve as reliable, domain-agnostic knowledge sources. Motivated by these advances, we explore their potential in a source-free setting and propose SPAR (Spatial Prior-Aware Regularization), a framework that uses class-agnostic binary masks derived from a frozen segmentation model [36] to regularize the detector’s feature space. SPAR encourages the student’s channel-mean activations to align with these masks, enhancing foreground structure and promoting object-focused responses. While SPAR improves the spatial quality of learned features, pseudo-labels in detection remain highly imbalanced, with abundant background regions and scarce positive samples. To stabilize training under this imbalance and mitigate residual class-label noise, we introduce IRPL (Imbalance-aware Noise Robust Pseudo-Labeling). It is a classification loss designed to promote balanced and noise-tolerant learning by down-weighting overconfident teacher-student agreements, reinforcing underrepresented categories, and countering the inherent background dominance in detection. Together, SPAR and IRPL complement each other: SPAR strengthens the structural integrity of object representations, while IRPL ensures balanced and robust supervision. We further provide a theoretical analysis linking these contributions to tighter localization and classification error bounds. Overall, our framework advances object-focused adaptation under domain shift. Our main contributions are summarized as follows:
@@ 2 Related Works
UDAOD adapts a source-trained detector to an unlabeled target-domain when source data is available during adaptation [12, 31, 22, 10, 15]. Recent works like DINO Teacher [13] leverage foundation models by training a source-domain labeller with a frozen DINOv2 and then aligning the student’s patch features to a frozen DINO encoder. While this approach enriches supervision, it remains computationally demanding as the DINO encoder is used in an online manner for target adaptation. Our SPAR loss provides a contrasting design: it uses class-agnostic binary masks from a frozen open-vocabulary segmenter [36] that are computed once before adaptation, so the foundation model is not queried during training or inference and the added cost is a one-time preprocessing step.
@@
SFOD eliminates the need to access source data during adaptation [30, 20, 7, 29, 11]. IRG [30] refines pseudo-labels via instance-relation graphs, PETS [20] stabilizes teacher-student training with periodic exchanges, and Simple-SFOD [7] shows that careful self-training design can outperform complex architectures. SF-YOLO [29] adapts the YOLO detector in a source-free manner using target domain-specific augmentations and a teacher-student communication mechanism. DRU [11] dynamically retrains and updates the mean teacher to enhance stability and performance during adaptation. While these works focus on refining pseudo-labels or stabilizing training, they overlook the importance of object-focused feature representation. Our proposed model addresses both the problems in a lightweight (simple-by-design, zero inference overhead), effective manner.
@@
Noisy-label learning introduces objectives that suppress the impact of corrupted labels [32]. However, a direct adaptation of  [32] fails in SFOD (detailed in Section 5.2) due to proposal imbalance, heavy class skew, and confirmation bias in teacher-student agreements. We extend it with foreground/background weighting, and entropy regularization, which together make IRPL effective for detection.
@@ 3.2 FALCON-SFOD
Motivation and Overview. As observed in [30, 20, 7, 11], a prevalent problem in mean-teacher self-training is that pseudo class-labels tend to be noisy due to domain shifts. Beyond this well-known issue, we identify that domain shift weakens the detector’s object-focus in SFOD: spatial activations for true objects are diluted by background clutter, which degrades localization and cascades into misclassification (see Fig. 1). We propose FALCON-SFOD, coupling two dedicated objectives within the standard Mean-Teacher framework: SPAR for spatial focus regularization and IRPL for pseudo class-label robustness. These losses are integrated with the conventional localization objective to enhance robustness during source-free adaptation.
Spatial Prior-Aware Regularization (SPAR). Consider the detector $h(x)=f(g(x))$ (Sec. 3.1). The feature extractor outputs an activation tensor $a=g(x)\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ are its height, width, and channel count. Taking the channel-wise mean yields $A\in\mathbb{R}^{H\times W}$, which highlights spatial locations with high average activations and thus usually traces foreground objects [23]. Under a domain shift, however, this map becomes contaminated by background clutter: true objects may fade while background areas are spuriously accentuated (as illustrated in Fig. 1). We derive a foreground prior once, offline, by running a frozen open-vocabulary segmentation model on target images. We discard the class information and binarize these masks, where pixels belonging to any segmented region are set to $1$ (foreground) and the rest to 0 (background). The prior is not queried during training or inference. For a target image $x_{i}^{t}$, let $A_{G}(x_{i}^{t})\in[0,1]^{H^{\prime}\times W^{\prime}}$ be the prior mask and $A_{S}(x_{i}^{t})\in[0,1]^{H\times W}$ be the student channel-mean map rescaled to $H^{\prime}\times W^{\prime}$ to match the prior. We encourage agreement using a mean $\ell_{1}$ term plus Dice::
@@
Imbalance-aware Noise Robust Pseudo-Labeling (IRPL). Most state-of-the-art SFOD methods minimize cross-entropy loss [30, 20, 7], which has an unbounded gradient that allows a single corrupted pseudo class-label to dominate training. Additionally, object detection has an inherent foreground-background imbalance which affects the stability of the student training. We therefore introduce IRPL, which mitigates this negative impact of noisy pseudo class-labels by adaptively recalibrating the per-box classification loss and effectively re-weights foreground-vs-background in the loss. For student probabilities $f_{c}^{st}\!\in\!\mathbb{R}^{K+1}$, let $\mathbf{p}=f_{c}^{st}$ and $t=\arg\max_{k}p_{k}$. The following transform rescales the peak while making sure that elements of $\mathbf{p}^{\prime}$ sum up to 1.
@@ 4 Theoretical Insights
Lemma 1 employs the class-conditional transition matrix $T\!\in\![0,1]^{(K+1)\times(K+1)}$, where $T_{ji}=\Pr[\hat{c}=i\mid c=j]$ represents the label noise induced by the mean-teacher: an exponential-moving-average (EMA) copy of the student whose deterministic predictions form a stochastic channel over the data distribution [27]. The diagonal element $T_{jj}$ is the teacher’s per-class hit-rate, and we define $\lambda=\min_{j}T_{jj}>0$ to rule out classes the teacher never recognizes, a minimal condition for identifiability of the clean risk under arbitrary label noise [21]. Lemma 1 therefore shows that mean-teacher asymmetry costs only a multiplicative factor $1/\lambda$; when the teacher is perfect ($\lambda=1$) the bound reduces to the standard clean-risk expression. Next, we show in Lemma 2 the regression risk under noisy pseudo bounding box labels.
@@
UDAOD SSAL [22] (NeurIPS’21) 45.1 47.4 59.4 24.5 50 25.7 26 38.7 39.6 51.8 45.6
PT [1] (ICML’22) 40.2 48.8 59.7 30.7 51.8 30.6 35.4 44.5 42.7 55.1 60.2
MTM [33] (AAAI’24) 51 53.4 67.2 37.2 54.4 41.6 38.4 47.7 48.9 58.1 -
SEEN-DA [15] (CVPR’25) 58.5 64.5 71.7 42 61.2 54.8 47.1 59.9 57.5 66.8 67.1
SFOD SFOD [17] (AAAI’21) 21.7 44.0 40.4 32.2 11.8 25.3 34.5 34.3 30.6 42.3 43.6
SFOD-Mosaic [18] (AAAI’21) 25.5 44.5 40.7 33.2 22.2 28.4 34.1 39.0 33.5 42.9 44.6
LODS [16] (CVPR’22) 34.0 45.7 48.8 27.3 39.7 19.6 33.2 37.8 35.8 - 43.9
IRG [30] (CVPR’23) 37.4 45.2 51.9 24.4 39.6 25.2 31.5 41.6 37.1 45.2 46.9
PETS [20] (ICCV’23) 42.0 48.7 56.3 19.3 39.3 5.5 34.2 41.6 35.9 57.8 47.0
Simple-SFOD [7] (ECCV’24) 40.9 48 58.9 29.6 51.9 50.2 36.2 44.1 45.0 55.4 46.2
SF-YOLO [29] (ECCV’24 w) 47.4 50.6 64.1 26.0 49.7 32.5 25.9 44.1 42.5 57.7 49.4
DRU [11] (ECCV’24) 48.3 51.5 62.5 26.2 43.2 34.1 34.2 48.6 43.7 58.7 45.1
@@ 5 Experiments
Datasets and Metrics. We use five publicly available datasets covering four domain shift scenarios. Cityscapes [3] is an urban street scene dataset comprising 5,000 finely annotated images collected from diverse cities and seasons, from which we use 2,925 images for training and 500 for validation. It includes eight categories: person, rider, car, truck, bus, train, motorcycle, and bicycle. Foggy Cityscapes [26] extends Cityscapes by overlaying synthetic fog at three intensity levels (0.005, 0.01, and 0.02) to simulate poor visibility conditions. KITTI [5] is an autonomous driving dataset consisting of 7,481 real-world street scene training images. Sim10k [9] provides 10,000 synthetic urban scene images of cars, rendered from the video game Grand Theft Auto. BDD100k [35] is a large-scale dataset comprising 100,000 driving scene images captured across various weather conditions. Following existing works, we report mean average precision (mAP) at 0.5 IoU threshold. To demonstrate the efficacy of our method on extreme domain shifts, we use following datasets: PascalVOC [4], COCO [19], FLIR [28], and Clipart [8].
Implementation Details. Our student model is trained using SGD with learning rates of 0.04 for source-only training and 0.0025 for target-domain adaptation, and a batch size of 4. The teacher model is updated via EMA with decay factor $\delta$ = 0.9996, using a confidence threshold 0.8. We use  [36] for getting the class-agnostic binary foreground masks for SPAR loss. The training parameters are identical for every dataset. All experiments are run on one NVIDIA RTX A6000 GPU. Results in the paper use [7] as the base model unless specified otherwise. Similar to  [30, 7], we use 0.02 fog-level images from the Foggy Cityscapes for adaptation.
@@ 5.1 Comparisons with State-of-the-art
Adaptation to Adverse Weather. To evaluate the adverse weather domain shifts, we perform adaptation from Cityscapes to Foggy Cityscapes. Table 1 shows that our method achieves a state-of-the-art (sota) mAP of 46.9%, outperforming DRU [11] by $3.2\%$, SF-YOLO [29] by $4.9\%$, and Simple-SFOD [7] by $1.9\%$. Especially, our method achieves notable improvements of 4.1%, 2.9%, and 2.4% in the under-represented and challenging classes train, bus, and motorcycle, respectively, compared to the previous sota in those categories.
@@
Small-scale to Large-scale. We select Cityscapes as the source-domain and BDD100k as the target-domain to study this shift. Following [20, 7], we focus on the seven categories shared with Cityscapes. As shown in Table 2, our method achieves an mAP of 36.9%, beating Simple-SFOD [7] by 2.6%.
@@
UDAOD SWDA [25] (CVPR’19) 15.2 45.7 29.5 30.2 17.1 21.2 18.4 25.3
CR-DA-Det [34] (CVPR’20) 19.5 46.3 31.3 31.4 17.3 23.8 18.9 26.9
MTM [33] (AAAI’24) 53.7 35.1 68.8 23.0 28.8 23.8 28.0 37.3
SFOD SFOD [17] (AAAI’21) 20.4 48.8 32.4 31.0 15.0 24.3 21.3 27.6
SFOD-Mosaic [18] (AAAI’21) 20.6 50.4 32.6 32.4 18.9 25.0 23.4 29.0
A2SFOD [2] (AAAI’23) 26.6 50.2 36.3 33.2 22.5 28.2 24.4 31.6
IRG [30] (CVPR’23) 31.4 59.7 32.8 39.9 16.7 26.9 21.5 32.7
PETS [20] (ICCV’23) 19.3 62.4 34.5 42.6 17.0 26.3 16.9 31.3
Simple-SFOD [7] (ECCV’24) 32 60 33.4 40.2 19.7 29.9 24.9 34.3
DRU [11] (ECCV’24) 27.1 62.7 36.9 45.8 22.7 32.5 28.1 36.6